In 2018, David, age 66, had adjusted gross income of $27,000. During the year he paid the following medical expenses: Prescription medicines $275Doctors $1350 Health insurance premiums insurance $400 Two pairs of eyeglasses $185 What amount can David deduct as medical expenses (after the adjusted gross income limitation) in calculating his itemized deductions for 2018? A $2710 b $2025 c $185 d $0
Answer #1

Answer:

Correct answer is:

c $185

Explanation:

For 2018, qualified medical expenses that exceed 7.5% of AGI is deductible.

Qualified medical expenses:

Prescription medicines = $275

Doctors = $1350

Health insurance premiums insurance = $400

Two pairs of eyeglasses = $185

Total = 275 +1350 + 400 + 185 = $2,210

7.5% of AGI = 27000 * 7.5% = $2,025

Deductible medical expenses (after the adjusted gross income limitation) = 2210 - 2025 = $185

Hence option c is correct and other options a, b, and d are incorrect.